Find the measure of 0 in the triangle. Do not use a calculator 7 3 0 5 O B =
Answer #1

we know the formula AC = AB² + BC² - 2.AB. BC. COS LB (7) = (3) + (5) - 2.3.5 . Cos LB > 49 = 9+25 - 30.COS LB > 30 COS LB 34This is the answe. Angle B= 120°